OpenStudy (anonymous):

5/2a-4/is greater than or equal to 15 solve the equation

OpenStudy (anonymous):

this is an absolute value inequality btw

OpenStudy (anonymous):

a>=(38)/(5) Does that help?

jhonyy9 (jhonyy9):

@tejadab268 step by step please !!! thank you

OpenStudy (anonymous):

@jhonyy9 (5)/(2)*a-4>=15 Multiply (5)/(2) by a to get (5a)/(2). (5a)/(2)-4>=15 Since -4 does not contain the variable to solve for, move it to the right-hand side of the inequality by adding 4 to both sides. (5a)/(2)>=4+15 Add 15 to 4 to get 19. (5a)/(2)>=19 Multiply each term in the inequality by 2. (5a)/(2)*2>=19*2 Simplify the left-hand side of the inequality by canceling the common factors. 5a>=19*2 Multiply 19 by 2 to get 38. 5a>=38 Divide each term in the inequality by 5. (5a)/(5)>=(38)/(5) Simplify the left-hand side of the inequality by canceling the common factors. a>=(38)/(5)